OpenStudy (anonymous):

Determine the domain of the function. f(x) = 4/(x^2)

OpenStudy (anonymous):

I think it's all real numbers except 0 but I'm not sure

OpenStudy (lyrae):

correct :)

OpenStudy (anonymous):

Could you help me with another?

OpenStudy (lyrae):

Sure

OpenStudy (anonymous):

domain of: \[f(x)=\sqrt{7+x}\]

OpenStudy (lyrae):

Do you have an idea what it might be?

OpenStudy (anonymous):

I don't remember how to solve the ones with square roots

OpenStudy (lyrae):

Okay, so for what numbers does a square root produce real answer?

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

@Lyrae

OpenStudy (lyrae):

Well, it's all real numbres grater then or eqal to 0 (try squaring a negative number on your calculator and you'll get an error or weird answer)

OpenStudy (anonymous):

okay

OpenStudy (lyrae):

So this gives us that what's inside that root has to be >= 0 This results in the following inequality: 7 + x >= 0

OpenStudy (lyrae):

Where x is all real numbres

OpenStudy (lyrae):

Do you know the final answer?

OpenStudy (anonymous):

So the domain is all real numbers?

OpenStudy (lyrae):

If you solve the inequality you get x >= -7 x is all real numbers, so the answer is: all numbres grater than or equal to -7 or in a more formal way \[\left\{x \in \mathbb{R} : x \ge -7\right\}\]

OpenStudy (anonymous):

Thank you sosossosososo much!!!!

OpenStudy (lyrae):

all real numbres* NP happy to help :)